FIRST-PRINCIPLES QUANTUM-CHEMICAL SIMULATIONS OF THE HYDRATED ELECTRON

\underline{JOHN M. HERBERT}, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

I will discuss our group’s ongoing efforts to elucidate the structure and dynamics of hydrated-electron clusters, (H_{2}O)_{n}^{−}, using ab initio quantum chemistry. Theoretical techniques include first-principles molecular dynamics and Monte Carlo simulations, as well as detailed analysis of the wave function associated with the “excess” electron. We apply these methods to predict and interpret both vibrational and photoelectron spectra of (H_{2}O)_{n}^{−}.